OpenStudy (anonymous):

FAN AND MEDAL! Which is a counterexample to the conjecture? Any number that is divisible by 3 is divisible by 9.

OpenStudy (anonymous):

6

OpenStudy (cwrw238):

think of a number less than 9 which is divisible by 3

OpenStudy (cwrw238):

12 is another counterexample

OpenStudy (anonymous):

It doesn't have to be less than 9, it just can't be a multiple of 9 :)

OpenStudy (anonymous):

so what is right?

OpenStudy (anonymous):

6 or 12??

OpenStudy (cwrw238):

both

OpenStudy (cwrw238):

i didnt say it had to be less than 9 - i was just giving a hint

OpenStudy (freckles):

any number with the prime factorization will work \[3 \cdot 2^k \cdot 5^k \cdot 7^k \cdot 11^{k}\cdot 13^k \cdot 17^k \cdots p^k \cdots \] where p is a prime number not equal to 3 and k is a non-negative integer

OpenStudy (freckles):

well and all those k's could be different :p

OpenStudy (freckles):

\[3 \cdot 2^{k_1} \cdot 5^{k_2} \cdot 7^{k_3} \cdot 11^{{k_4}}\cdot 13^{k_5} \cdot 17^{k_6} \cdots p^{k_n} \cdots\] where the k_i values are non-negative